# Chinese Academy of Sciences, Beijing, China

### v1

@@ Abstract
Large Vision Language Models have achieved fine-grained object perception, but the limitation of image resolution remains a significant obstacle to surpass the performance of task-specific experts in complex and dense scenarios. Such limitation further restricts the model’s potential to achieve nuanced visual and language referring in domains such as GUI Agents, Counting and etc. To address this issue, we introduce a unified high-resolution generalist model, Griffon v2, enabling flexible object referring with visual and textual prompts. To efficiently scaling up image resolution, we design a simple and lightweight down-sampling projector to overcome the input tokens constraint in Large Language Models. This design inherently preserves the complete contexts and fine details, and significantly improves multimodal perception ability especially for small objects. Building upon this, we further equip the model with visual-language co-referring capabilities through a plug-and-play visual tokenizer. It enables user-friendly interaction with flexible target images, free-form texts and even coordinates. Experiments demonstrate that Griffon v2 can localize any objects of interest with visual and textual referring, achieve state-of-the-art performance on REC, phrase grounding, and REG tasks, and outperform expert models in object detection and object counting. Data, codes and models will be released at https://github.com/jefferyZhan/Griffon.
Keywords:
Large Vision Language Models Multimodal perception High-resolution structure Visual-language co-referring
@@ 1 Introduction
Large Vision Language Models (LVLMs) [26, 2, 59, 13] show promising performance in region-level tasks like Referring Expression Comprehension (REC) after the breakthrough in image-text understanding [3, 12] and reasoning [16]. In particular, models like Griffon [55] have demonstrated more compelling perception capability in object detection task. This has further spurred the development of flexible references of objects beyond only textual descriptions for better user interaction.
@@
Figure 1: Overview of Griffon v2. Griffon v2 enables high-resolution input and visual-language co-referring for multimodal models. In terms of referring, a user can refer to an object through coordinates, textual descriptions, screenshotting and cross-image modes. Griffon v2 excels in localizing arbitrary objects and generate descriptions with co-referring across a wide range of scenarios.
@@
To address this issue, recent works have explored resolution enhancement and flexible visual-language referring in LVLMs. On the one hand, previous methods [11, 10] adopt a progressive training approach to gradually improve the resolution. However, the maximum input length of Large Language Models (LLMs) imposes a constraint on achieving higher image resolutions. Additionally, some approaches [25, 22, 27] divide the image to smaller patches and encode them separately for zooming in. This division-based design proves sub-optimal multimodal perception ability, which losses the contexts and edge details of patches and increases computation complexity [22]. On the other hand, prior researches have predominantly studied various forms of references [8, 57, 5] to improve the understanding of specific image content based on low-resolution images (e.g. 224$\times$224 or 448$\times$448). However, these methods often excel at perceiving prominent, and image-level objects, but fall short in accurately localizing and describing fine-grained local objects. Additionally, singular visual or language prompts alone either lack conversational abilities or are constrained by linguistic descriptions [15], failing to provide a comprehensive user-friendly interactive experience.
@@
In this paper, we propose Griffon v2 to open the evil of high-resolution and endow it with locating any objects of interest with visual-language co-referring. Instead of partitioning the high-resolution image into smaller patches, we employ a high-resolution visual encoder to direct extract representations, and design a simple and lightweight down-sampling projector with strided convolution to compress the length of visual tokens. The compressed visual tokens are then trained to align with text tokens and fed into LLM for further fusion like LLaVA [26]. Compared to complex resampler structure[25] and multiple sub-images mode, the proposed pipeline is parameter-efficient and computationally concise. More importantly, we build a visual-language co-referring paradigm to enhance the model’s fine-grained perception of high-resolution inputs, greatly expanding the model’s applicability. It supports local cropped images, texts, and coordinates prompting, and outputs coordinates of target objects or corresponding text descriptions, providing various interactive abilities, thereby mitigating the conversational deficiencies of singular visual prompting and the potential expressive ambiguity of textual prompting. Finally, we collected 12M publicly available localization data for pre-training and 900K instruction data for fine-tuning. We achieve state-of-the-art results in REC task, phrase grounding task, Referring Expression Generation (REG) task. Notably, our model outperforms object detection and object counting expert models for the first time.
@@
    We propose a high-resolution multimodal perception model for better local understanding, which inherently supports resolution up to 1K without image division.
@@
    We introduce a visual-language co-referring structure, which broadens the model’s scope of application and provide various interaction modes.
@@
    We have conducted experiments on a wide range of localization-related tasks and demonstrated state-of-the-art performance on REC, phrase grounding and REG. We quantitatively surpass expert models in object detection task object counting qualitatively.
@@ 2 Related Work
2.1 Large Vision Language Models with Localization
With the significant advancements made in LVLMs [26, 2, 13, 1], object localization as an essential foundation visual task has been explored with LVLMs. One branch of these methods [49, 39, 58, 23] focuses on invoking the detection expert models through an API. When the LVLM requires detection, the LLM issues instructions to call the detection model and receives the results for further processing. This approach increases the complexity of the model and can not enjoy the unified structure with LVLMs. Another set of methods [8, 4, 51, 11] focuses on fine-grained single-object localization, i.e. REC. These methods transform REC task data into instruction data and encode the coordinates with different representation approaches, enabling the LVLM itself to possess localization capabilities. Until the appearance of Griffon [55], the object localization capacity of LVLM is extended to multi-object and arbitrary granularity. Our model further proposes a high-resolution structure that elevates the detection capability of LVLM beyond expert models, which is also proved more effective and efficient than existing division-based resolution-enhanced methods [22, 25].
@@ 2.2 Object Referring in Multimodal Models
Since the application of LLMs [41, 1] in large multimodal models, textual descriptions have become the most straightforward method for object referring. Users naturally utilize this method to ask questions without even paying special attention, such as "What is in front of this cat?" Despite its simplicity, it struggles to distinctly refer to a specific object among numerous similar items, leading to ambiguous referring that may yield inaccurate results. With the growing importance of region comprehension, it has sparked the exploration of spatial references in complex and dense scenes. Current works have proposed spatial referrals with textual coordinates [8, 4, 7], learnable embeddings [33], and Region of Interest features [57]. Some approaches [5, 51, 1] upgrade from the user-friendly perspective, supporting arrows, marks, and so on. Due to the limitations of their focus on tasks and granularity, these approaches primarily concentrate on referring within a single modality, also without further consideration for high resolution. In contrast, our model simultaneously considers perception, localization, and comprehension at different granularity. Therefore, we employ a visual-language co-referring approach to enable accurate references in various tasks while remaining user-friendly.
3 Approach
In this section, we start with a brief overview of our Griffon v2 as shown in Fig. 2. As the key foundation, we describe the high-resolution structure design in Sec. 3.2. Then, we represent the visual-language co-referring in Sec. 3.3 and the training pipeline in Sec. 3.4.
@@ 3.1 Overview
As depicted in Fig. 2, Griffon-V2 employs an auto-regression paradigm, seamlessly integrating referring and grounding tasks within a unified language modeling objective. For a given input image $X_{v}$, we leverage a visual encoder, adapted to high-resolution by bilinear interpolation from the pretrained EVA2-CLIP-L/14@336 [40] model to extract high-resolution visual features (e.g. over 1K). Simultaneously, the designed down-sampling projector transforms these image features into visual embedding tokens $H_{v}$. In cases where a user customizes a visual prompt $X_{q}$ with a screenshot or an additional target image, region embeddings $H_{q}$ are derived from the visual referring image using the visual tokenizer. Meanwhile, we inherently support language references, $X_{ins}$, to identify objects, such as textual descriptions and coordinates, which are projected into text embeddings as $H_{ins}$ after tokenization. Subsequently, the visual prompt or text embeddings together with image embeddings undergo processing using the LLM, specifically LLaMA2-13B [41], resulting in the generation of the desired answer $X_{a}$. Notably, the representation of bounding boxes eschews the use of special tokens and specialized heads and instead incorporate textual numerical values, thereby maintaining simplicity.
3.2 High-Resolution Scaling with Down-Sample Projector
[…]
To address these challenges, we introduces a high-resolution structure, incorporating a down-sampling projector capable of supporting a resolution more than 1K.
Figure 3: Performance and visual token number comparison of different resolution. We fix the stride at 2 with the kernel size at 3 for the convolution layer and vary the resolution for analysis. The 448-resolution model employs a two-layer linear projector [26], whereas the others utilize our down-sampling projector. our method is capable of extracting features while preserving details, thereby reducing token redundancy.
Upon receiving an input image of arbitrary dimensions, we resize it to a predefined input height and width, denoted as $X_{v}\in\mathbb{R}^{H\times W\times 3}$. Instead of employing image division, we encode the image using trainable high-resolution visual encoder adapted from pretrained model by bilinear interpolation, yielding visual features $Z_{v}=f_{V}(X_{v})$. In contrast to low-resolution fixed encoders, our approach excels in capturing finer details. As illustrated in the left part of Fig. 2, we subsequently introduce a lightweight down-sampling projector to abstract essential features with compression under the guidance of the training objective and connect visual features to word embedding space. Specifically, a strided convolution layer ($Conv$) is simply applied to down-sample the features by $S$, and a projection matrix $W$ is utilized to convert $Z_{v}$ into visual embedding tokens:
@@ 3.2 High-Resolution Scaling with Down-Sample Projector
It can be inferred that the final number of visual tokens is quadratically related to the input size and the convolution stride, i.e. $H,W$ and $S$. To demonstrate the effectiveness of our design, we employ the convolution layer with kernel size at 3 and stride at 2, and investigate several different resolutions in Fig. 3. As depicted in Fig. 3, comparing the results at resolutions of 700 and 448 which use two-layer linear projection [26] without reduction, using the resolution of 700 with our structure shows a higher mAP and fewer tokens. This indicates that our method is capable of extracting features while preserving details, thereby reducing token redundancy. It is also observed that further increasing the resolution can lead to improved performance.
@@ 3.3 Visual-Language Co-Referring
Figure 4: Examples of visual-language co-referring usage in conversations.
@@
Recognizing the limitations of singular visual or language referring demonstrated before, lack of conversational ability and potential ambiguity respectively, we capitalize on CLIP’s adeptness in extracting robust semantic features to facilitate visual-language co-referring. This approach mitigates these drawbacks and finds broader applicability under high-resolution conditions.
@@
Visual Referring. While various existing methods propose diverse visual reference forms, ranging from marking on the raw image to object masking, either of them alter the raw image content [1] or require specially designed extractor [51]. To strike a balance between efficiency and referring accuracy, we opt for a straightforward method, using visual referring image to represent the target of interest. Specifically, as depicted in the middle of Fig. 2, given a user-provided referring image $X_{q}$, we employ a region encoder, i.e. EVA-CLIP-L/14@224 [40] pretrained ViT to extract visul prompt features $Z_{q}$. To align the LLM’s understanding of visual references with textual references, we project the region features into the word embedding space, yielding the visual prompt token $H_{q}$. Based on the robust semantic representation capability of $[CLS]$ token, pre-trained on a dataset of billions of instances [36], we discard other tokens following previous applications and utilize the $[CLS]$ token to represent the specific object. This approach allows us to refer to the object at the category level, eliminating overly specific information from pixel patches and maintaining the stability of reference. To accommodate diverse conversation scenarios, we use “<region>” in instructions to represent the referred object as shown in the Fig. 4 (b). Before inputting into the LLM, we replace this placeholder with the embedding of the referred object, seamlessly integrating it for optimal coherence.
@@
Language Referring. Textual referring is an inherent capability of LLM-based multimodal models. Despite the simplicity of textual referring, it faces the challenge of resulting in wrong answer with ambiguous referring. This has naturally prompted us to consider whether associating texts with regions could allow LLMs to comprehend the corresponding object based on textual-region reference and thereby fulfill task instructions and keep the simplicity advantage. Therefore, we further support textual coordinate references as illustrated in the Fig. 4 (a). In additional to the detailed description, an object can also be referred by its top-left and bottom-right coordinates. In this case, it is processed the same as textual descriptions, involving tokenization and embedding into the word space as $H_{ins}$, depicted at the right side of Fig. 2. Through training with instruction-following data whose object references are transformed into textual coordinates, we demonstrate that LLMs can accurately refer to objects in complex and dense scenes based on textual coordinates of objects. Through collaborative referring involving both text and vision, our approach achieves optimal referring performance and flexible interaction.

### v2

# Griffon v2: Advancing Multimodal Perception with High-Resolution Scaling and Visual-Language Co-Referring

@@ Abstract
Large Vision Language Models have achieved fine-grained object perception, but the limitation of image resolution remains a significant obstacle to surpassing the performance of task-specific experts in complex and dense scenarios. Such limitation further restricts the model’s potential to achieve nuanced visual and language referring in domains such as GUI Agents, counting, etc. To address this issue, we introduce a unified high-resolution generalist model, Griffon v2, enabling flexible object referring with visual and textual prompts. To efficiently scale up image resolution, we design a simple and lightweight down-sampling projector to overcome the input tokens constraint in Large Language Models. This design inherently preserves the complete contexts and fine details and significantly improves multimodal perception ability, especially for small objects. Building upon this, we further equip the model with visual-language co-referring capabilities through a plug-and-play visual tokenizer. It enables user-friendly interaction with flexible target images, free-form texts, and even coordinates. Experiments demonstrate that Griffon v2 can localize objects of interest with visual and textual referring, achieve state-of-the-art performance on REC and phrase grounding, and outperform expert models in object detection, object counting, and REG. Data and codes are released at https://github.com/jefferyZhan/Griffon.
@@ 1 Introduction
Large Vision Language Models (LVLMs) [29, 2, 56, 15] show promising performance in region-level tasks like Referring Expression Comprehension (REC) after the breakthrough in image-text understanding [3, 11] and reasoning [19]. In particular, models like Griffon [53] have demonstrated more compelling perception capability in object detection tasks. This has further spurred the development of flexible references of objects beyond only textual descriptions for better user interaction.
Figure 1: Overview of Griffon v2. Griffon v2 enables high-resolution input and seamless visual-language co-referencing for LVLMs, allowing users to refer to objects through coordinates, textual descriptions, screenshots, or cross-image modes. It excels in localizing arbitrary objects and generating descriptions with precise co-referencing across diverse scenarios.
@@
To address this issue, recent works have explored resolution enhancement and flexible visual-language referring in LVLMs. On the one hand, previous methods [13, 12] adopt a progressive training approach to gradually improve the resolution. However, the maximum input length of Large Language Models (LLMs) imposes a constraint on achieving higher image resolutions. Additionally, some approaches [28, 25, 30] divide the image into smaller patches and encode them separately for zooming in. This division-based design proves sub-optimal multimodal perception ability, which loses the contexts and edge details of patches and increases computation complexity [25]. On the other hand, prior research has predominantly studied various forms of references [9, 54, 6] to improve the understanding of specific image content based on low-resolution images (e.g. 224 or 448). However, these methods often excel at perceiving prominent, and image-level objects, but fall short in accurately localizing and describing fine-grained local objects. Additionally, singular visual or language prompts alone either lack conversational abilities or are constrained by linguistic descriptions [17], failing to provide a comprehensive user-friendly interactive experience.
@@
In this paper, we propose Griffon v2 to unveil the direct high-resolution design and endow it with locating any objects of interest with visual-language co-referring. Instead of partitioning the high-resolution image into smaller patches, we employ a high-resolution visual encoder to directly extract representations, and design a simple and lightweight down-sampling projector with strided convolution to compress the length of visual tokens. The compressed visual tokens are then trained to align with text tokens and fed into LLM for further fusion like LLaVA [29]. Compared to complex resampler structure[28] and image partitioning methods mode[25, 30], the proposed direct high-resolution pipeline preserves context and advances dense instance-level tasks a lot with competitive VQA performances. It is also parameter-efficient and computationally concise. More importantly, we build a visual-language co-referring paradigm to enhance the model’s fine-grained perception of high-resolution inputs, greatly expanding the model’s applicability. It supports local cropped images, texts, and coordinates prompting, and outputs coordinates of target objects or corresponding text descriptions, providing various interactive abilities, thereby mitigating the conversational deficiencies of singular visual prompting and the potential expressive ambiguity of textual prompting. Finally, we collected 12M publicly available localization data for pre-training and 900K instruction data for fine-tuning. We achieve advanced results in the REC task, phrase grounding task, and Referring Expression Generation (REG) task. Notably, our model outperforms several object detection and object counting expert models for the first time.
@@
    We propose a high-resolution multimodal perception model for better local understanding and less global context loss, better suited for challenging tasks such as intensive multi-object detection and counting in complex scenarios.
@@
    We introduce a visual-language co-referring structure, which broadens the model’s scope of application and provides various interaction modes.
@@
    We have conducted experiments on a wide range of localization-related tasks and demonstrated state-of-the-art performance on REC, phrase grounding, and REG. We surpass expert models in object detection and object counting task quantitatively and qualitatively.
@@ 2 Related Work
2.1 LVLMs with Localization
With the significant advancements made in LVLMs[29, 2, 15, 1], object localization as an essential foundation visual task has been explored with LVLMs. One branch of these methods [49, 42, 55, 26] focuses on invoking the detection expert models through an API. When the LVLM requires detection, the LLM issues instructions to call the detection model and receives the results for further processing. This approach increases the complexity of the model and can not enjoy the unified structure with LVLMs. Another set of methods [9, 4, 50, 13] focuses on fine-grained single-object localization, i.e. REC. These methods transform REC task data into instruction data and encode the coordinates with different representation approaches, enabling the LVLM itself to possess localization capabilities. However, methods like Ferret [50] fall short in more complex tasks like object detection and more challenging scenarios with lots of small objects under the low-resolution. Until the appearance of Griffon[53], which supports coarser granularity object detection, the object localization capacity of LVLM is extended to multi-object and arbitrary granularity. Our model further proposes a high-resolution structure that elevates the detection and object counting capability of LVLM beyond expert models, which has also proved more effective and efficient than existing division-based resolution-enhanced methods [25, 28]. By enhancing the object detection and enabling object counting, our work address a broader spectrum of granularity in comprehension, perception and localization.
@@ 2.2 Object Referring in Multimodal Models
Since the application of LLMs [44, 1] in large multimodal models, textual descriptions have become the most straightforward method for object referring. Users naturally utilize this method to ask questions without even paying special attention, such as “What is in front of this cat?” Despite its simplicity, it struggles to distinctly refer to specific classes of objects that are difficult to describe in dense and complex scenes, such as the cell shown in Figure 1. The growing importance of region comprehension has sparked the exploration of spatial references in complex and dense scenes. Current works have proposed spatial referrals with textual coordinates [9, 4, 8], learnable embeddings [36], and Region of Interest features [54]. Some approaches [6, 50, 1] upgrade from the user-friendly perspective, supporting arrows, marks, and so on. Although these methods enhance the convenience of user interaction, they primarily focus on designing visual prompts to distinguish one specific object from others and are mostly applied in text-output VQA tasks. In contrast, our model uses visual referring to represent a class of objects, not limited to distinction of individual objects. Additionally, we introduce bounding box output to enable dense scenario understanding in LMMs for the first time. The proposed visual-language co-referring architecture empowers our model to handle various tasks while remaining user-friendly.
3 Methodology
In this section, we start with a brief overview of our Griffon v2. As the key foundation, we first describe the high-resolution structure design. Then, we represent the visual-language co-referring and the training pipeline.
@@ 3.1 Overview
As depicted in Figure 2(a), Griffon v2 employs an auto-regression paradigm, seamlessly integrating referring and grounding tasks within a unified language modeling objective. For a given input image $X_{v}$, we leverage a visual encoder, adapted to high-resolution by bilinear interpolation from the pre-trained EVA2-CLIP-L/14@336 [43] model to extract high-resolution visual features (e.g. over 1K). Simultaneously, the designed down-sampling projector transforms these image features into visual embedding tokens $H_{v}$. In cases where a user customizes a visual prompt $X_{q}$ with a screenshot or an additional target image, region embeddings $H_{q}$ are derived from the visual referring image using the visual tokenizer. Meanwhile, we inherently support language references, $X_{ins}$, to identify objects, such as textual descriptions and coordinates, which are projected into text embeddings as $H_{ins}$ after tokenization. Subsequently, the visual prompt or text embeddings together with image embeddings undergo processing using the LLM, specifically Llama2-13B [44], resulting in the generation of the desired answer $X_{a}$. Notably, the representation of bounding boxes eschews the use of special tokens and specialized heads and instead incorporates textual numerical values, thereby maintaining simplicity.
3.2 Efficient High-Resolution Scaling Design
High resolution has been empirically proven to enhance the accurate interpretation of nuanced visual features, thereby benefiting both visual tasks and vision-language tasks. In previous methods [25], images are commonly partitioned into smaller patches, enlarging their total resolution to 762 or 896. However, this approach faces limitations in terms of contextual coherence and the representation of edge details among patches. While progressive scaling methods can extend resolution to 784 (patch size==14) within the constraint of 4096 input tokens, they are unsuitable for handling outputs involving extensive long texts and significantly increase the computation complexity.
To address these challenges, we introduce a high-resolution structure, incorporating a down-sampling projector capable of supporting a resolution of more than 1K. Upon receiving an input image of arbitrary dimensions, we resize it to a predefined input height and width, denoted as $X_{v}\in{R}^{H\times W\times 3}$. Instead of employing image division, we encode the image using a trainable high-resolution visual encoder adapted from a pre-trained model by bilinear interpolation, yielding visual features $Z_{v}=f_{V}(X_{v})$. In contrast to low-resolution fixed encoders, our approach excels in capturing finer details. As illustrated in the left part of Figure 2(a), we subsequently introduce a lightweight down-sampling projector to abstract essential features with compression under the guidance of the training objective and connect visual features to word embedding space. Specifically, a strided convolution layer ($Conv$) is simply applied to down-sample the features by $S$, and a projection matrix $W$ is utilized to convert $Z_{v}$ into visual embedding tokens:
@@ 3.2 Efficient High-Resolution Scaling Design
It can be inferred that the final number of visual tokens is quadratically related to the input size and the convolution stride, i.e. $H,W$, and $S$. To demonstrate the effectiveness of our design, we employ the convolution layer with kernel size at 3 and stride at 2 and investigate several different resolutions in Figure 3. As depicted in Figure 3, comparing the results at resolutions of 700 and 448 which use two-layer linear projection [29] without reduction, using the resolution of 700 with our structure shows a higher mAP and fewer tokens. This indicates that our method is capable of extracting features while preserving details, thereby reducing token redundancy. It is also observed that further increasing the resolution can lead to improved performance.
Figure 3: Performance and visual token number comparison of different resolutions under the same training data. The 448-resolution model, i.e. Griffon v1, employs a two-layer linear projector, whereas the others utilize our down-sampling projector (Stride==2, Kernel Size==3). Our method is capable of extracting features while preserving details, thereby reducing token redundancy.
@@ 3.3 Visual-Language Co-Referring
Recognizing the limitations of singular visual or language referring demonstrated before, lack of conversational ability, and potential ambiguity respectively, we capitalize on CLIP’s adeptness in extracting robust semantic features to facilitate visual-language co-referring. This approach mitigates these drawbacks and finds broader applicability under high-resolution conditions.
@@
Visual Referring. While various existing methods propose diverse visual reference forms, ranging from marking on the raw image to object masking, either of them alters the raw image content [1] or requires a specially designed extractor [50]. To strike a balance between efficiency and referring accuracy, we opt for a straightforward method, using visual referring images to represent the target of interest. Specifically, as depicted in the middle of Figure2(a), given a user-provided referring image $X_{q}$, we employ a region encoder, i.e. EVA-CLIP-L/14@224 [43] pre-trained ViT to extract visual prompt features $Z_{q}$. To align the LLM’s understanding of visual references with textual references, we project the region features into the word embedding space, yielding the visual prompt token $H_{q}$. Based on the robust semantic representation capability of $[CLS]$ token, pre-trained on a dataset of billions of instances [39], we discard other tokens following previous applications [39, 43] and utilize the $[CLS]$ token to represent the specific object. This approach allows us to refer to the object at the category level, eliminating overly specific information from pixel patches and maintaining the stability of reference. To accommodate diverse conversation scenarios, we use “<<region>>” in instructions to represent the referred object as shown in Figure 2(b). Before inputting into the LLM, we replace this placeholder with the embedding of the referred object, seamlessly integrating it for optimal coherence.
@@
Language Referring. Textual referring is an inherent capability of LLM-based multimodal models. Despite the simplicity of textual referring, it faces the challenge of resulting in wrong answers with ambiguous referring. This has naturally prompted us to consider whether associating texts with regions could allow LLMs to comprehend the corresponding object based on textual-region reference and thereby fulfill task instructions and keep the simplicity advantage. Therefore, we further support textual coordinate references as illustrated in Figure 2(b). In addition to the detailed description, an object can also be referred to by its top-left and bottom-right coordinates. In this case, it is processed the same as textual descriptions, involving tokenization and embedding into the word space as $H_{ins}$, depicted at the right side of Figure 2(a). Through training with instruction-following data whose object references are transformed into textual coordinates, we demonstrate that LLMs can accurately refer to objects in complex and dense scenes based on the textual coordinates of objects. Through collaborative referring involving both text and vision, our approach achieves optimal referring performance and flexible interaction.